A company estimates that 0.2% of their products will fail after the original warranty period but within 2 years of the purchase, with a replacement cost of $150.

If they offer a 2 year extended warranty for $9, what is the company's expected value of each warranty sold?
1
Expert's answer
2020-06-29T19:21:49-0400


Solution:

let x be the outcomes are $9 and -($150-9)=-$141


p(x) be the probabilities of outcomes are 1-0.002=0.998 and 0.002 respectively


Expected value ="\Sigma x*p(x)"

=(9*0.998)+(-141*0.002)

=8.982-0.282

=8.7


Answer:$8.7